OpenStudy (anonymous):

David's company reimburses his expenses on food, lodging, and conveyance during business trips. The company pays $60 a day for food and lodging and $0.65 for each mile traveled. David drove 600 miles and was reimbursed $3390. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did David spend on this trip? (1 points)

OpenStudy (anonymous):

please help !

OpenStudy (anonymous):

3390=x60+(0.65)600 3390=x60+390 subtraction property of equality 3390-390=x60+390-390 3000=x60 divide x=50

OpenStudy (anonymous):

@Todd_The_Best im still a little confused , so for part a the equation is 3390=x60+(0.65)600 and then part b would be 3390=x60+390 what do i do now ?

OpenStudy (anonymous):

Oh, I see. For part C all I did was multiply 0.65 times 600 and subtracted it from 3390. Then I divided that by 60 and got the answer

OpenStudy (anonymous):

i got it ! thank you @mipie13

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

Thanx Mippie

OpenStudy (anonymous):

Can you help me with one more? @Todd_The_Best @mipie13

OpenStudy (anonymous):

Mike poured 15 cups of water into a rectangular container measuring 10 inches by 6 inches by 10 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.

OpenStudy (anonymous):

@mipie13 @Todd_The_Best

OpenStudy (anonymous):

first find the volume of the container, so 10 times 6 times 3.9

OpenStudy (anonymous):

10 times 6 times 3.9 is 234 @lucyv7129

OpenStudy (anonymous):

so that is the answer to part a

OpenStudy (anonymous):

Okay got part a

OpenStudy (anonymous):

for part b, simply multiply 15 cups by 14.4, because 1 cup is equal to 14.4 cubic inches

OpenStudy (anonymous):

whats the next step for part b

OpenStudy (anonymous):

okay so that would be 216

OpenStudy (anonymous):

@lucyv7129

OpenStudy (anonymous):

yep, do you know how to do part c, try it

OpenStudy (anonymous):

could someone help me please ? Mike poured 15 cups of water into a rectangular container measuring 10 inches by 6 inches by 10 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.